If SSR = 23 and SST = 92, compute the coefficient of determination, r?, and interpret its meaning. 12 (Type an integer or a dWill rate, thank you in advance.
Answer #1

Solution :

From given that ,

r 2 = SSR / SST = 23 / 92 = 0.25

C. Pothe variation in the independentvarable can be explained by the varation in the dependent variable.